Required information

[The following information applies to the questions displayed below.]

Beavis Construction Company was the low bidder on a construction project to build an earthen dam for $1,750,000. The project was begun in 2017 and completed in 2018. Cost and other data are presented below:

2017 2018
Costs incurred during the year $ 469,000 $ 1,050,000
Estimated costs to complete 871,000 0
Billings during the year 480,000 1,270,000
Cash collections during the year 380,000 1,425,000

Assume that Beavis recognizes revenue on this contract over time according to percentage of completion.

Required:
Compute the amount of gross profit recognized during 2017 and 2018.

Amount of gross profit recognized in 2017 ____

Amount of gross profit recognized in 2018 ____
Answer #1

Gross profit = [(costs to date / estimated total costs * expected profit) - previously recognized profits.

Costs to date in 2017 = 469,000

Estimated total costs in 2017 = 469,000 + 871,000 = 1,340,000

Expected profit in 2017 = 1,750,000 - 1,340,000 = 410,000

Previously recognized profit in 2017 = 0

Gross profit in 2017 = [(469,000 / 1,340,000 * 410,000) - 0] = 143,500

Costs to date in 2018 = 469,000 + 1,050,000 = 1,519,000

Estimated total costs in 2018 = 1,519,000

Expected profit in 2018 = 231,000

Previously recognized profit in 2017 = 143,500

Gross profit in 2018 = [(1,519,000 / 1,519,000 * 231,000) - 143,500] = 87,500